a balloon is 50m high its angle of elevation from observer A is 45 and from observer B it is 30 what is the distance between observer b and observer a from the baloon


1
Expert's answer
2022-06-06T23:18:03-0400
"x=\dfrac{h}{\tan\theta}"

"x_A=\dfrac{50}{\tan45\degree}=50(m)"

"x_B=\dfrac{50}{\tan30\degree}=50\sqrt{3}(m)"

Maximum distance


"d_{max}=x_A+x_B=50+50\sqrt{3}=50(1+\sqrt{3})(m)"

"\approx136.6(m)"

Minimum distance


"d_{max}=-x_A+x_B=-50+50\sqrt{3}=50(\sqrt{3}-1)(m)"

"\approx36.6(m)"